# Many-Body Localization in the Age of Classical Computing

### v2

@@ IV.3.1 Single parameter scaling
Let us reconsider the quantity $\overline{X}$ that indicates the break-down of ergodicity in the system, and examine its dependence on the system size $L$ and disorder strength $W$, denoted as $\overline{X}=\overline{X}(L,W)$.
A large body of numerical works starting from [460], through the more accurate [393] and continuing to these days, treated the ergodicity breaking transition of MBL as a ‘usual’ phase transition, i.e., as if it can be described by a simple fixed point of a renormalization group analysis [5, 105] with a single relevant parameter. We will review this concept, assumptions, and results in the following but we anticipate that the results are unsatisfactory. In particular, the critical exponents which best fit the data are in contradiction with general bounds (to be discussed below), and the irrelevant corrections are large, with a considerable flow of the apparent fixed point with system size [^10] A possible exception is given by certain Floquet models as argued in [596], see Sec. X.4 .
In a single-parameter scaling hypothesis the results for $\overline{X}(L,W)$ sufficiently close to the transition (i.e., when all the irrelevant parameters flow can be neglected) can be interpreted as depending solely only on the product of a scaling variable $\chi(w)$ and the system size $L$
@@
where $f_{0}$ is a certain scaling function, $w\equiv(W-W_{C})/W_{C}$ is the dimensionless disorder, and the scaling variable, generally written as $\chi(w)=\sum_{n}a_{n}w^{n}$, can be assumed to be proportional to $w$, $\chi(w)\sim w$. The critical exponent $\nu$ is given by the linearization of the beta function, see Eq. (13), of a relevant variable close to the critical point $w=0$. We remind the reader that if one considers $X$ as an RG “charge" (like a coupling or a simple function of couplings in the Hamiltonian), one can use the RG flow to organize different finite size systems on the same orbit [105]. This is done by means of a beta function
$\beta\equiv\frac{d\ln X}{d\ln L}.$ (13)
The one-parameter scaling gives rise to the hypothesis that $\beta$ is solely a function of $X$, i.e., $\beta=\beta(X)$. Assuming a simple critical fixed point with the behavior $\beta(X)\simeq s(X-X_{0})$ close to the fixed point, then integrating the equation we have $\ln(|X(L)-X_{0}|/|X(L_{0})-X_{0}|)=X_{0}s\ln(L/L_{0})$. So $|X(L)-X_{0}|\simeq|X(L_{0})-X_{0}|(L/L_{0})^{1/\nu}$, with $\nu=1/(sX_{0})$ the critical exponent. Since at finite $L_{0}$ everything is analytic in $w$, we must have $|X(L_{0})-X_{0}|\propto\chi(w),$ where $\chi(w)$ is an analytic function of $w$. In the generic case we expect $\chi(w)\simeq w+O(w^{2})$, and therefore we can write $X(L)=X_{0}+c(L/\xi)^{1/\nu}$, where $\xi$ is a correlation length that exhibits a power-law divergence
@@
However, it turns out that applying the single parameter scaling hypothesis (12)-(14) to the results at the putative ETH-MBL transition presents significant difficulties. Under a natural assumption that the function $f_{0}(x)$, defined in (12), is continuous and monotonous in the vicinity of the transition, the single parameter scaling hypothesis by construction fails to reproduce the characteristic, significant drift of the crossing point observed in results at the ETH-MBL crossover. Indeed, in this scenario, the equality $\overline{X}(L_{1},W^{*})=\overline{X}(L_{2},W^{*})$ holding for two system sizes $L_{1}\neq L_{2}$ at the crossing point $W^{*}$ implies, through (12)-(14), that $W^{*}=W_{C}$. Consequently, while $W^{*}=W_{C}$ for the behavior schematically represented in Fig. 2(a), the single parameter scaling is not well suited for the ETH-MBL crossover, c.f. Fig. 2(b). Nevertheless, if one disregards the drift of the crossing point $W^{*}$, for instance by restricting the results to a few largest system sizes available, the finite size scaling ansatz (12)-(14) can yield a visually appealing data collapses at the ETH-MBL crossover. Unfortunately, the resulting critical exponent $\nu$ obtained in this approach is close to $1$ [460, 326, 393], in apparent contradiction with the Harris-Chayes criterion $\nu\geq 2$ [265, 119] applicable to 1D systems [115, 316].
Recent phenomenological approaches [242, 177] propose that the ETH-MBL transition may be of Berezinskii–Kosterlitz–Thouless (BKT) [337], suggesting to replace (14) with the following behavior of the correlation length
@@
The BKT scaling (15) is characterized by a much slower scaling of the results with system size when compared to the power-law scenario (14). To demonstrate this, let us consider fixing a specific value $x_{0}$ of our ergodicity breaking indicator, such that the equality $\overline{X}(L,W)=x_{0}$ fixes a parametric dependence of $W=W_{x_{0}}(L)$. Assuming that the system undergoes a transition at which $\overline{X}$ exhibits a step-like behavior in the $L\to\infty$ limit, we observe that for $x_{0}$ smaller (larger) than the value of $\overline{X}$ corresponding to the critical point, $W_{x_{0}}(L)$ converges with increasing $L$ to the critical disorder strength $W_{C}$ from below (above), compare Fig. 2. For the power-law scenario, from (12) and (14), we find that
@@
Since the understanding of this subject is still evolving, it is important to mention that for the Anderson transition on random graphs (which is not in the universality class of any finite-dimensional Anderson transition), it has been pointed out in [671] (see also [595, 216]) that the one-parameter scaling theory of [5] needs to be amended. The RG equations of [5] on infinite dimensional lattices are inherently different from the ones on finite-dimensional lattices. Among other things, this implies that the critical point is localized itself (the level statistics is Poisson, and the critical eigenstates are localized [657]). Since the localized fixed point and the critical point cannot be distinguished, the linearization of the beta function close to the critical point that should give rise to the critical exponent $\nu$ and to the usual one-parameter scaling behavior does not work tout-court. When one considers the whole beta function (13), together with the now non-negligible irrelevant parameters flow at the critical point, the whole picture can be mapped to the BKT transition described before. It could therefore be the case that the MBL transition is in the same universality class as the Anderson transition in infinite dimensions (something already implied by the work [22] and reiterated in [156, 20, 671]). Work to establish this connection, following [671, 23], and to re-analyze the existing results from this new point of view, is in progress.
@@ IV.3.2 Sub-leading corrections to the scaling
where $f_{0}(wL^{1/\nu})=F(wL^{1/\nu},0)$ and $f_{1}(wL^{1/\nu})=\partial_{x}F(wL^{1/\nu},x)|_{x=0}$. Comparing (19) with (12), we see that the irrelevant variable yields a correction to the scaling through the term $\psi L^{y}f_{1}(wL^{1/\nu})$. In contrast to the single parameter scaling form (12), the scaling ansatz (19) describes a drift of the crossing point $W^{*}$ with increasing system size $L$. The relation between $W^{*}$ and $L$ is determined by the exponents $\nu$ and $y$, as well as the functional forms of the functions $f_{0}$ and $f_{1}$; see [595] for an example of Anderson transition on random graphs. Thus, the scaling ansatz (19) is more suited for numerical results at the ETH-MBL crossover as exemplified in [596], which predicted $\nu\approx 2$ for 1D interacting disordered Floquet system, in consistence with the Harris-Chayes criterion. However, the scaling form (19) contains multiple free parameters while the numerical results for many-body systems are available only in a limited range of system sizes, which severely restricts the predictive power of a finite scaling analysis with (19). This is illustrated by the scaling analysis of Anderson localization transition on random regular graphs. While the exact knowledge of the critical disorder strength makes the analysis more constrained, the presently available numerical data alone are insufficient to clearly determine the value of the critical exponent $\nu$, see [486, 595, 657, 215, 216, 78].
@@
which leads to a correlation length that depends not only on the disorder strength $W$, for instance through (14) or (15), but also on the system size $L$, $\chi=\chi(L,W)$. Different scenarios for the transition can be described by different scaling of $\widetilde{W}_{C}(L)$ with the system size $L$. Conventional scenario assumes that the transition occurs at a finite critical disorder strength in the thermodynamic limit $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}W_{C}$. However, Eq. (20) also introduces possibilities of unconventional scaling, in which the critical point diverges in a certain fashion with system size, $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$, or that the regime at small $W$ shrinks to a single point in the large system size limit $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}0$.
@@ IV.3.3 Unconventional scaling limits
Since the introduction of the scaling theory of localization [5], it was understood that the 2D Anderson model remains localized at any disorder strength $W$. Nevertheless, early numerical studies of the model, e.g. [185, 372], found a regime of disorder strengths at which eigenstates of the model appeared to be delocalized. In contrast, a diagrammatic approach of [685], relevant for a weak disorder, predicts the breakdown of transport due to interference corrections at a lengthscale which, for asymptotically small $W$, must behave like $\xi\propto e^{a/W^{2}}$, where $a$ is a constant. The lengthscale $\xi$ is identified, within the self-consistent theory of localization, with the localization length of the 2D Anderson model [359]. This is due to the fact that the $\beta$ function has a double zero rather than a simple zero at the critical point. This is the way the RG captures the effects of divergent series of weak-localization corrections. Numerical analysis of 2D Anderson model, relying on a direct extraction of the localization length [403], or, recently, on a study of the average gap ratio $\overline{r}$ [693], suggests a single-parameter scaling with the length scale $\xi^{\prime}\propto e^{a^{\prime}/W^{\mu^{\prime}}}$, with $\mu^{\prime}\approx 1$ for the accessible system sizes. While a gradual adherence of the numerical results to the perturbative prediction (which becomes exact as $W\to 0$) can be observed [357], the Anderson model in 2D is an important reference point for studies of the localization transitions. It shows that data collapses over even seemingly wide intervals of $L$ may yield the spurious value of the exponent $\mu^{\prime}=1$. The delocalized-localized crossover can be interpreted, via (20), as an unconventional transition which is getting sharper with increasing $L$, but eventually the delocalized phase shrinks to a point, $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}0$. In some sense, explained by the work of [671], the Anderson model in 2D shares a similarity with the Anderson model on random regular graphs, since in both cases the critical point coincides with one of the “gaussian fixed points" (localized or ergodic). In 2D it is the ergodic fixed point, in the random regular graphs it is the localized fixed point. In both cases larger than power-law critical lengths emerge since a ‘naive’ analysis based on a simple zero of the beta function must fail. Both cases need to be treated using a more complex scaling analysis, and this is a cautionary tale for the studies of MBL as well.
@@
Disordered spin-1/2 chains with power-law interactions provide a further example in which an unconventional scaling limit allows for interesting ergodicity breaking phenomena [656, 240]. The presence of power-law interactions characterized by an exponent $a$ leads to an absence of MBL phase in that model, $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$, when the thermodynamic limit is taken in the usual manner, i.e. by keeping the tunneling matrix elements, disorder and interaction strengths fixed. Identifying the $L$ dependence of the transition point $\widetilde{W}_{C}(L)$ on the combination of the exponent $a$ and the dimensionality of the system $d$ yields a variable $w=(W-\widetilde{W}_{C}(L))/\widetilde{W}_{C}(L)$. Ref. [656] suggests that the variable $w$ allows for tuning the disordered spin-1/2 chains with power-law interactions across an MBL transition in the $L\to\infty$ limit, see also Sec. X.5. Similarly, changing $w$ leads to sharp transition between ergodic and MBL phases in the thermodynamic limit in the avalanche model considered by [240].

### v3

@@ IV.3.1 Single parameter scaling
Let us reconsider the quantity $\overline{X}$ that indicates the breakdown of ergodicity in the system, and examine its dependence on the system size $L$ and disorder strength $W$, denoted as $\overline{X}=\overline{X}(L,W)$.
A large body of numerical works starting from [498], through the more accurate [430] and continuing to these days, treated the ergodicity breaking transition of MBL as a ‘usual’ phase transition, i.e., as if it can be described by a simple fixed point of a renormalization group analysis [5, 108] with a single relevant parameter. We will review this concept, assumptions, and results in the following but we anticipate that the results are unsatisfactory. In particular, the critical exponents which best fit the data are in contradiction with general bounds (to be discussed below), and the irrelevant corrections are usually large, with a considerable flow of the apparent fixed point with system size.
In a single-parameter scaling hypothesis the results for $\overline{X}(L,W)$ sufficiently close to the transition (i.e., when all the irrelevant parameter flows can be neglected) can be interpreted as depending solely only on the product of a scaling variable $\chi(w)$ and the system size $L$,
@@
where $f_{0}$ is a certain scaling function, $w\equiv(W-W_{C})/W_{C}$ is the dimensionless disorder, and the scaling variable, generally written as $\chi(w)=\sum_{n}a_{n}w^{n}$, can be assumed to be proportional to $w$, $\chi(w)\sim w$. The critical exponent $\nu$ is given by the linearization of the beta function, see Eq. (13), of a relevant variable close to the critical point $w=0$. We remind the reader that if one considers $\overline{X}$ as an renormalization group “charge" (like a coupling or a simple function of couplings in the Hamiltonian), one can use the RG flow to organize different finite size systems on the same orbit [108]. This is done by means of a beta function
$\beta\equiv\frac{d\ln\overline{X}}{d\ln L}.$ (13)
The one-parameter scaling gives rise to the hypothesis that $\beta$ is solely a function of $\overline{X}$, i.e., $\beta=\beta(\overline{X})$. Assuming a simple critical fixed point with the behavior $\beta(\overline{X})\simeq s(\overline{X}-\overline{X}_{0})$ close to the fixed point, then integrating the equation we have $\ln(|\overline{X}(L)-\overline{X}_{0}|/|\overline{X}(L_{0})-\overline{X}_{0}|)=\overline{X}_{0}s\ln(L/L_{0})$. Hence, $|\overline{X}(L)-\overline{X}_{0}|\simeq|X(L_{0})-X_{0}|(L/L_{0})^{1/\nu}$, with $\nu=1/(s\overline{X}_{0})$ the critical exponent. Since at finite $L_{0}$ everything is analytic in $w$, we must have $|\overline{X}(L_{0})-\overline{X}_{0}|\propto\chi(w),$ where $\chi(w)$ is an analytic function of $w$. In the generic case we expect $\chi(w)\simeq w+O(w^{2})$, and therefore we can write $\overline{X}(L)=\overline{X}_{0}+c(L/\xi)^{1/\nu}$, where $\xi$ is a correlation length that exhibits a power-law divergence,
@@
However, it turns out that applying the single parameter scaling hypothesis (12)-(14) to the results at the putative ETH-MBL transition presents significant difficulties. Under a natural assumption that the function $f_{0}(x)$, defined in (12), is continuous and monotonous in the vicinity of the transition, the single parameter scaling hypothesis by construction fails to reproduce the characteristic, significant drift of the crossing point observed in results at the ETH-MBL crossover. Indeed, in this scenario, the equality $\overline{X}(L_{1},W^{*})=\overline{X}(L_{2},W^{*})$ holding for two system sizes $L_{1}\neq L_{2}$ at the crossing point $W^{*}$, implies, through (12)-(14), that $W^{*}=W_{C}$. Consequently, while $W^{*}=W_{C}$ for the behavior schematically represented in Fig. 2(a), the single parameter scaling is not well suited for the ETH-MBL crossover, c.f. Fig. 2(b). Nevertheless, if one disregards the drift of the crossing point $W^{*}$, for instance by restricting the results to a few largest system sizes available, the finite size scaling ansatz (12)-(14) can yield a visually appealing data collapses at the ETH-MBL crossover. Unfortunately, the resulting critical exponent $\nu$ obtained in this approach is close to $1$ [498, 353, 430], in apparent contradiction with the Harris-Chayes criterion $\nu\geq 2$ [284, 125] applicable to 1D systems [119, 342].
Several phenomenological approaches [258, 187] proposed that the ETH-MBL transition may be of the Berezinskii–Kosterlitz–Thouless (BKT) origin [366], suggesting to replace (14) with the following behavior of the correlation length,
@@
The BKT scaling (15) is characterized by a much slower scaling of the results with system size when compared to the power-law scenario (14). To demonstrate this, let us consider fixing a specific value $x_{0}$ of our ergodicity breaking indicator, and define a system size dependent disorder strength $W_{x_{0}}(L)$ such that $\overline{X}(L,W_{x_{0}}(L))=x_{0}$. Assuming that the system undergoes a transition at which $\overline{X}$ exhibits a step-like behavior in the $L\to\infty$ limit, we observe that for $x_{0}$ smaller (larger) than the value of $\overline{X}$ corresponding to the critical point, $W_{x_{0}}(L)$ converges with increasing $L$ to the critical disorder strength $W_{C}$ from above (below), compare Fig. 2. For the power-law scenario, from (12) and (14), we find that
@@
Since the understanding of this subject is still evolving, it is important to mention that for the Anderson transition on random graphs (which is not in the universality class of any finite-dimensional Anderson transition), it has been pointed out in [725] (see also [644, 231]) that the one-parameter scaling theory of [5] needs to be amended. The RG equations of [5] on infinite dimensional lattices are inherently different from the ones on finite-dimensional lattices. Among other things, this implies that the critical point is localized itself (the level statistics is Poisson, and the critical eigenstates are localized [709]). Since the localized fixed point and the critical point cannot be distinguished, the linearization of the beta function close to the critical point that should give rise to the critical exponent $\nu$ and to the usual one-parameter scaling behavior does not work tout-court. When one considers the whole beta function (13), together with the now non-negligible irrelevant parameters flow at the critical point, the whole picture can be mapped to the BKT transition described before. It could therefore be the case that the MBL transition is in the same universality class as the Anderson transition in infinite dimensions (something already implied by the [24] and reiterated in [163, 22, 725]). Work to establish this connection, following [725, 25], and to re-analyze the existing results from this new point of view, is in progress.
@@ IV.3.2 Sub-leading corrections to the scaling
where $f_{0}(wL^{1/\nu})=F(wL^{1/\nu},0)$ and $f_{1}(wL^{1/\nu})=\partial_{x}F(wL^{1/\nu},x)|_{x=0}$. Comparing (19) with (12), we see that the irrelevant variable yields a correction to the scaling through the term $\psi L^{y}f_{1}(wL^{1/\nu})$. In contrast to the single parameter scaling form (12), the scaling ansatz (19) describes a drift of the crossing point $W^{*}$ with increasing system size $L$. The relation between $W^{*}$ and $L$ is determined by the exponents $\nu$ and $y$, as well as the functional forms of the functions $f_{0}$ and $f_{1}$; see [644] for an example of Anderson transition on random graphs. Thus, the scaling ansatz (19) is more suited for numerical results at the ETH-MBL crossover as exemplified in [645], which predicted $\nu\approx 2$ for 1D interacting disordered Floquet system, in consistence with the Harris-Chayes criterion. However, the scaling form (19) contains multiple free parameters while the numerical results for many-body systems are available only in a limited range of system sizes, which severely restricts the predictive power of a finite scaling analysis with (19). This is illustrated by the scaling analysis of Anderson localization transition on random regular graphs. While the exact knowledge of the critical disorder strength makes the analysis more constrained, the presently available numerical results alone are insufficient to clearly determine the value of the critical exponent $\nu$, see [525, 644, 709, 230, 231, 81].
@@
which leads to a correlation length that depends not only on the disorder strength $W$, for instance through (14) or (15), but also on the system size $L$, $\chi=\chi(L,W)$. Different scenarios for the transition can be described by different scaling of $\widetilde{W}_{C}(L)$ with the system size $L$. Conventional scenario assumes that the transition occurs at a finite critical disorder strength in the thermodynamic limit $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}W_{C}$. However, Eq. (20) also introduces possibilities of unconventional scaling, in which the critical point diverges in a certain fashion with system size, $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$, or that the regime at small $W$ shrinks to a single point in the large system size limit, $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}0$.
@@ IV.3.3 Unconventional scaling limits
Since the introduction of the scaling theory of localization [5], it was understood that the 2D Anderson model remains localized at any disorder strength $W$. Nevertheless, early numerical studies of the model, e.g., [195, 406], found a regime of disorder strengths at which eigenstates of the model appeared to be delocalized. In contrast, a diagrammatic approach of [740], relevant for a weak disorder, predicts the breakdown of transport due to interference corrections at a lengthscale which, for asymptotically small $W$, must behave like $\xi\propto e^{a/W^{2}}$, where $a$ is a constant. The lengthscale $\xi$ is identified, within the self-consistent theory of localization, with the localization length of the 2D Anderson model [389]. This is due to the fact that the $\beta$ function has a double zero rather than a simple zero at the critical point [^18] At simple zero $\beta(\overline{X})=0$ and $d\beta(\overline{X})/d\overline{X}\neq 0$. At double zero $\beta(\overline{X})=0$ and $d\beta(\overline{X})/d\overline{X}=0$.. This is the way the RG captures the effects of divergent series of weak-localization corrections. Numerical analysis of 2D Anderson model, relying on a direct extraction of the localization length [439], or, recently, on a study of the average gap ratio $\overline{r}$ [749], suggests a single-parameter scaling with the length scale $\xi^{\prime}\propto e^{a^{\prime}/W^{\mu^{\prime}}}$, with $\mu^{\prime}\approx 1$ for the accessible system sizes. While a gradual adherence of the numerical results to the perturbative prediction (which becomes exact as $W\to 0$) can be observed [387], the Anderson model in 2D is an important reference point for studies of the localization transitions. It shows that data collapses over even seemingly wide intervals of $L$ may yield the unconventional value of the exponent $\mu^{\prime}=1$. The delocalized-localized crossover can be interpreted, via (20), as an unconventional transition which is getting sharper with increasing $L$, but eventually the delocalized phase shrinks to a point, $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}0$. In some sense, explained by the work of [725], the Anderson model in 2D shares a similarity with the Anderson model on random regular graphs, since in both cases the critical point coincides with one of the “gaussian fixed points" (localized or ergodic). In 2D it is the ergodic fixed point [^19] For 2D Anderson model, at the ergodic fixed point the fractal dimension $D_{1}$ (see Sec. VI.2.1) is equal to unity, and the $\beta$ function can be expanded around the $D_{1}=1$ fixed point as $\beta(D_{1})=-(1-D_{1})^{2}$ [25]., in the random regular graphs it is the localized fixed point. In both cases larger than power-law critical lengths emerge since a ‘naive’ analysis based on a simple zero of the beta function must fail. Both cases need to be treated using a more complex scaling analysis, and this is a cautionary tale for the studies of MBL as well.
Quantity Measures Ergodic regime MBL regime Section
Average gap ratio $\overline{r}$ short-range $\overline{r}\stackrel{{\scriptstyle L\gg 1}}{{\longrightarrow}}\overline{r}_{\mathrm{GOE}}=0.5307(1)$ $\overline{r}\stackrel{{\scriptstyle L\gg 1}}{{\longrightarrow}}\overline{r}_{\mathrm{PS}}=0.386(1)$ VI.1.1
level statistics
Spectral form factor $K(t)$ long-range $K(t)=K_{\mathrm{GOE}}(t)$ for $t>t_{\mathrm{Th}}$ $K(t)=K_{\mathrm{PS}}(t)=1$, $t_{\mathrm{Th}}\geq t_{\mathrm{H}}$ VI.1.2
level statistics
Participation entropy $S_{q}$ wave function $S_{q}=\log_{2}(\mathcal{N})+c_{q}$, $c_{q}<0$ $S_{q}=D_{q}\log_{2}(\mathcal{N})+c_{q}$, $D_{q}<1$ VI.2.1
delocalization
Entanglement entropy $S_{A}$ eigenstate $S_{A}=S_{RMT}\propto L_{A}$ (volume-law) $S_{A}=\mathrm{const}$ (area-law) VI.2.2
entanglement
Spectral function $f^{2}(\omega)$ off-diagonal $f^{2}(\omega)=\mathrm{const}$ for $\omega<\omega_{\mathrm{Th}}$ $f^{2}(\omega)=C_{0}\delta(\omega)$$+$$f^{2}_{\mathrm{reg}}(\omega)$, $C_{0}$$>$00 VI.3.1
matrix elements
Fidelity susceptibility $\chi_{\mathrm{typ}}$ eigenstate $\chi_{\mathrm{typ}}\sim\mathcal{N}$ $\chi_{\mathrm{typ}}/\mathcal{N}\stackrel{{\scriptstyle L\gg 1}}{{\longrightarrow}}0$ VI.3.2
sensitivity
Density correlation $C(t)$ relaxation $C(t)\stackrel{{\scriptstyle t\to\infty}}{{\longrightarrow}}0$ for $L\gg 1$ $C(t)\stackrel{{\scriptstyle t\to\infty}}{{\longrightarrow}}C_{0}>0$ for $L\gg 1$ VII.1
dynamics
Configurational entangle- entanglement $S_{c}(t)\propto t$ $S_{c}(t)\propto\log_{2}(t)$ VII.2.1
ment entropy $S_{c}(t)$ generation
Number entropy $S_{n}(t)$ particle number $S_{n}(t)\propto\log_{2}(t)$ $S_{n}(t)\propto\mathrm{const}$ VII.2.2
fluctuations
Nested commutator operator $||\mathcal{L}^{n}\hat{A}||\propto(n/\ln(n))^{n}$ $||\mathcal{L}^{n}\hat{A}||\leq(\mathrm{const})^{n}$ VII.3.4
norm $||\mathcal{L}^{n}\hat{A}||$ spreading
Spin conductivity $\sigma(\omega)$ spin $\sigma(\omega)\stackrel{{\scriptstyle\omega\ll 1}}{{=}}\sigma_{0}+\zeta\omega^{1-2/z}$, $z\geq 2$ $\sigma(\omega)\stackrel{{\scriptstyle\omega\ll 1}}{{=}}0$ VIII.1.1
transport
Two-point density density $\Delta x(t)\propto t^{1/z}$ with $z\geq 2$ $\Delta x(t)=\mathrm{const}$ VIII.1.2
correlator $\Phi(x,t)$ spreading
Spin current $j$ spin $j\propto L^{1-z}$ with $z\geq 2$ $j\propto\exp(-\gamma L)$ with $\gamma>0$ VIII.1.3
transport
Table 1: Summary of the quantities employed as ergodicity breaking indicators and their behavior in ergodic and MBL regimes, together with the number of the Section in which they are defined and considered. Here and throughout, we consider systems defined on one-dimensional lattices consisting of $L$ sites, $L_{A}$ is size of subsystem $A$ which defines a bipartition of the system, $t$ is time (measured in tunneling times of the system), $\mathcal{N}$ is the dimension of the relevant sector of the many-body Hilbert space, $t_{\mathrm{Th}}\,(\omega_{\mathrm{Th}})$ is the Thouless time (energy), $t_{\mathrm{H}}$ ($\omega_{\mathrm{H}}$) is the Heisenberg time (energy), $z$ is the dynamical critical exponent. The persistence of the behavior associated with the ergodic (MBL) regime in the asymptotic limits $t\to\infty$ and $L\to\infty$ over an extended range of disorder strengths $W$ indicates the existence of an ergodic (MBL) phase.
@@
Disordered spin-1/2 chains with power-law interactions provide a further example in which an unconventional scaling limit allows for interesting ergodicity breaking phenomena [708, 256]. The presence of power-law interactions characterized by an exponent $a$ leads to an absence of MBL phase in that model, $\widetilde{W}_{C}(L)\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$, when the thermodynamic limit is taken in the usual manner, i.e., by keeping the tunneling matrix elements, disorder and interaction strengths fixed. The thermodynamic limit can be taken in unconventional way. Ref. [708] observes that the variable $w=(W-\widetilde{W}_{C}(L))/\widetilde{W}_{C}(L)$ allows for tuning the disordered spin-1/2 chains with power-law interactions across a transition in the $L\to\infty$ limit, see also Sec. X.5. Similarly, changing $w$ leads to sharp transition between ergodic and MBL phases in the thermodynamic limit in the avalanche model considered by [256].